OpenStudy (anonymous):

antiderivative of sqrt of x(x+2)dx

OpenStudy (anonymous):

\[\int\limits \sqrt{x}(x+2)\]

OpenStudy (anonymous):

start by multiplying using exponential notation and get \[\int (x^{\frac{3}{2}}+2x^{\frac{1}{2}})dx\] and then use the power rule backwards \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

Thank you